# Meta-Learning an In-Context Transformer Model of Human Higher Visual Cortex

### v1

@@ Abstract
Understanding functional representations within higher visual cortex is a fundamental question in computational neuroscience. While artificial neural networks pretrained on large-scale datasets exhibit striking representational alignment with human neural responses, learning image-computable models of visual cortex relies on individual-level, large-scale fMRI datasets. The necessity for expensive, time-intensive, and often impractical data acquisition limits the generalizability of encoders to new subjects and stimuli. BraInCoRL  uses in-context learning to predict voxelwise neural responses from few-shot examples without any additional finetuning for novel subjects and stimuli. We leverage a transformer architecture that can flexibly condition on a variable number of in-context image stimuli, learning an inductive bias over multiple subjects. During training, we explicitly optimize the model for in-context learning. By jointly conditioning on image features and voxel activations, our model learns to directly generate better performing voxelwise models of higher visual cortex. We demonstrate that BraInCoRL consistently outperforms existing voxelwise encoder designs in a low-data regime when evaluated on entirely novel images, while also exhibiting strong test-time scaling behavior. The model also generalizes to an entirely new visual fMRI dataset, which uses different subjects and fMRI data acquisition parameters. Further, BraInCoRL facilitates better interpretability of neural signals in higher visual cortex by attending to semantically relevant stimuli. Finally, we show that our framework enables interpretable mappings from natural language queries to voxel selectivity.
@@ 1 Introduction
Human visual cortex transforms raw sensory input into behaviorally-relevant representations of the world. While early visual areas are characterized by retinotopic organization and selective tuning to simple features such as edges and orientation gradients [1, 2, 3, 4, 5], higher-order visual areas demonstrate selectivity to more abstract semantics and categories. While this functional organization is largely consistent across individuals at a coarse scale, the spatial distribution and fine-grained semantic selectivity within visual cortex varies due to structural differences, developmental experience, and life-long learning [6, 7, 8, 9, 10, 11]. Such functional inter-subject differences pose a fundamental challenge in constructing generalizable models of higher visual cortex that can adapt to subject-specific neural organization without exhaustive data collection for every individual.
Recent advances in deep learning offer a promising avenue for addressing this challenge. Vision models pretrained on large-scale image datasets not only achieve strong object recognition performance, but also recapitulate hierarchical processing patterns observed in biological vision [12, 13, 14]. While these models may encapsulate some universal principles of visual processing [15], they do not inherently account for individual differences in cortical organization. To close the gap between artificial and biological systems, researchers have developed image-computable fMRI encoders – models that predict brain activity from visual stimuli  [16]. These encoders typically regress image features onto voxelwise brain responses using subject-specific data, acting as computational probes of visual processing. Unfortunately, current approaches require many hours of costly fMRI scans per subject to fit these mappings – a prohibitive bottleneck for scalability to new populations, stimuli, and tasks, especially in clinical settings where collecting large amounts of data is difficult.
@@
Figure 1: BraInCoRL: Meta-Learning an In-Context Visual Cortex Encoder. (a) The voxelwise brain encoding problem. For each voxel, there is a response function that maps from visual stimuli to voxel activation. In practice, we can only observe the noisy measurements from fMRI. The goal is to infer an image-computable function for each voxel to predict its activation. (b) BraInCoRL treats each voxel as a meta-learning task, and samples (image, response) pairs from multiple subjects. During testing, the model is conditioned on a small number of novel images and measurements from a new subject and directly outputs the function parameters. (c) From left to right, the explained variance from the full model trained on 9,000 images from one subject, BraInCoRL with only $100$ in-context images from the new subject, and a baseline ridge regression also with $100$ images (for this baseline, voxelwise regularization is determined using 5-way cross-validation). Our method achieves much higher data efficiency than baseline. (d) Explained variance as a function of in-context support set size. As the in-context support set size increases from 0 to 1,000, BraInCoRL steadily improves and approaches the fully trained reference model fit to converge on each subject’s full 9,000-image training set, demonstrating high prediction accuracy and data efficiency.
@@ 2 Related work
Computational modeling of neural data often involves two complementary approaches: encoding models that map from stimuli to neural activations, and decoding models that map from neural data to stimuli [16, 17, 18, 19, 20, 21, 22, 23, 24]. The development of both approaches has been facilitated by advances in machine learning models. For encoding models, the dominant approach is to combine pretrained deep feature extractors with linear voxelwise weights [25, 26, 27, 28, 29, 30]. More recent approaches have proposed to leverage transformers [31, 32] to learn the relationship between brain regions of a single subject. Most similar to our framework is the pioneering work by Adeli et al. [31] and Beliy & Wasserman et al. [33] which uses an auto-decoder based transformer network for multi-subject voxelwise encoding; However these approaches still require fine-tuning for novel subjects. More generally, encoders have been used to investigate the coding properties in higher-order visual areas [34, 35, 36, 37, 38, 39, 40, 41]. Encoders have been further combined with image-generation models  [42, 43, 44, 45, 46, 47, 48, 49] or language-generation models [50, 51] to explore semantic selectivity. Recent progress on large generative models has enabled stimulus reconstruction from fMRI, EEG, and MEG signals for images [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63], videos [64, 65, 66, 67, 68, 69, 70], and language/audio [71, 72, 73, 74, 75, 76, 77].
@@
Human visual cortex exhibits a hierarchical organization from primary visual to higher-order visual areas. The higher visual cortex is characterized by a tiling of semantically specialization. Approaches using functional localizers have identified category selective regions in higher visual that are responsive to faces [78, 79, 80, 81], places [82, 83], bodies [84], objects [85, 86], food [87, 88, 89], and words [90, 91]. While the spatial location of these broad category-selective regions are generally consistent across subjects [92], significant inter-individual variability exists in their anatomical location, spatial extent, and response profiles [6, 7, 8, 9, 93, 94, 95, 96, 97]. Accurately characterizing visual processing in higher-order visual areas necessitates subject-specific encoding models that capture individual diversity [98].
@@
Our work builds upon meta-learning and in-context learning (ICL). Meta-learning trains models to "learn to learn" from a distribution of tasks, enabling quick adaptation to new tasks with few examples, via methods like meta-optimization [99, 100, 101] or metric-based approaches [102]. More recently, ICL has emerged as a powerful capability in Large Language Models [103, 104], where models adapt to new tasks at inference time solely based on examples provided in their prompt without any parameter updates [105, 106]. This has led to hypotheses that ICL is an emergent form of implicit meta-learning, where transformers effectively learn underlying learning algorithms during pre-training [107, 108]. Our goal is to learn the structure of functions that map between visual stimuli and voxelwise brain response. Our framework combines meta-training (across voxels and subjects) and in-context learning (across stimuli) to enable training free adaptation to novel subjects.
@@ 3.2 Meta-Learning an In-Context Transformer
For a single voxel we define a support set of $p$ images and neural responses $\{(x_{1},\beta_{1}),(x_{2},\beta_{2}),...,(x_{p},\beta_{p})\}$, where $x_{i}\in\mathbb{R}^{m}$ is the image embedding vector extracted by a frozen image feature extractor $\phi(\cdot)$, i.e., $x_{i}=\phi(I_{i})$, and $\beta_{i}\in\mathbb{R}$ is the voxel’s response observed for image $I_{i}$. Each pair is concatenated to form context tokens $c_{i}=[x_{i};v_{i}]$, and the full context is defined as $\{c_{1},\dots,c_{p}\}$. Unlike traditional in-context inference in large language models, where there is a query concatenated to the end of the context, we avoid expensive per-voxel inference by directly generating the parameters for the voxelwise visual response function. During training, we optimize the BraInCoRL transformer $T$ with parameters $\theta$ such that it outputs voxel response function $f$ with parameters $\omega$:
@@ 3.3 Test-time Context Scaling
At test time, when we encounter a new subject, we assume we have access to a small set of novel images and the corresponding brain responses – we want to predict a voxelwise encoder. While our voxelwise parameterization successfully resolves the challenge of unknown test-time anatomy and geometry, the challenge of unknown test-time context size remains. Unlike transformers in language models, where the output is dependent on the order of the samples, we want our model to be explicitly invariant to input order. In order to facilitate variable length context, we utilize logit scaling [109, 110, 111]. Assuming a query/key ($q,k$) with $d_{k}$ features and a length $c$ context:
$\alpha_{\text{orig}}=\frac{q\cdot k}{\sqrt{d_{k}}};\quad\alpha_{\text{scaled}}=\frac{\log{(c)}\cdot q\cdot k}{\sqrt{d_{k}}}$ (4)
We find this method effectively enables context scaling when trained with variable length context. While the hypernetwork could, in principle, parameterize any class of neural encoders (e.g., MLPs, convolution, attention layers), prior studies utilizing brain data have largely used linear parameterizations that map from deep network features to voxel responses [13, 14], and find that such a choice offers high performance and interpretability. Given features $x\in\mathbb{R}^{1\times q}$ from a pretrained neural network $x=\phi(\mathcal{I})$, we adopt the same setup and predict the final voxel response:
@@ 4 Experiments
Figure 3: Evaluation on NSD. (a) Prediction explained variance of BraInCoRL improves on novel subjects with larger in-context support set size, outperforming within-subject ridge regression and approaching the fully trained reference model fit on each subject’s full 9,000-image training set, using far less data. (b) Ablation (100 support images) comparing BraInCoRL variants: the original model trained while holding out the novel subject’s 9,000 test-time support images (“HO”), a BraInCoRL model trained without this holdout (“no HO”), and a pretraining-only BraInCoRL model, alongside the within-subject ridge baseline. Results show that finetuning with real fMRI data improves performance, and holding out the test subject’s image data does not hinder generalization. (c) Voxelwise explained variance from BraInCoRL (100 images) is strongly correlated with fully trained reference models across different visual encoder backbones.
Table 1: Voxelwise performance across five category-selective regions. Explained variance is shown for our in-context model (“BraInCoRL”) that uses just 100 in-context images, the fully trained reference model fit to converge on each subject’s full 9,000-image training set (“Fully Trained”), and within-subject ridge regression baselines (100, 300 within-subject test images), plus the FsAverage map averages over other subjects. Our model outperforms both subject-wise and anatomical baselines, and demonstrates strong data-efficiency.
@@ 4.1 Setup
We primarily perform experiments with the Natural Scenes Dataset (NSD) [112], but then validate with the BOLD5000 dataset [113]. Both are large-scale neural datasets: NSD is the largest 7T fMRI image viewing dataset available, where eight subjects each viewed $\sim 10,000$ images; BOLD5000 is a 3T dataset, where four subjects each viewed $\sim 5000$ images. In NSD each image was viewed up to three times, while in BOLD5000 only a subset of images were viewed up to four times. For NSD, of the eight subjects, four subjects (S1, S2, S5, S7) completed scanning and are the focus of our analysis in the main paper. The results of other subjects are presented in the supplemental materials. For each subject, $\sim 9,000$ images are unique to each other, while $\sim 1,000$ are viewed by all eight subjects. The NSD stimuli were sourced from the MS COCO dataset (as were a subset of the BOLD5000 stimuli). Unless otherwise noted, we perform our analysis in subject native volume space (func1pt8mm) for NSD, where the voxelwise betas are z-scored within each session then averaged across repeats of the same stimulus. In order to rigorously evaluate BraInCoRL for a given subject, we use the 3 × 9,000 unique images viewed by the other three subjects as the meta-training data. During the ROI-wise evaluation for NSD, we follow prior work [48] and apply a $t$-statistic cutoff of $t>2$ by leveraging independent functional localizer data provided with the dataset to threshold the originally broad definitions. During quantitative evaluation, we follow prior work [114] and apply a voxel quality cutoff of $\texttt{ncsnr}>0.2$. For BOLD5000, we use a model trained on four NSD subjects (S1, S2, S5, S7). Following the suggestion of BOLD5000 authors, we only model stimuli with 4 repeats and apply a cutoff of $\texttt{ncsnr}>0.3$. Voxel responses are averaged across repeats.

### v2

@@ Abstract
Understanding functional representations within higher visual cortex is a fundamental question in computational neuroscience. While artificial neural networks pretrained on large-scale datasets exhibit striking representational alignment with human neural responses, learning image-computable models of visual cortex relies on individual-level, large-scale fMRI datasets. The necessity for expensive, time-intensive, and often impractical data acquisition limits the generalizability of encoders to new subjects and stimuli. BraInCoRL uses in-context learning to predict voxelwise neural responses from few-shot examples without any additional finetuning for novel subjects and stimuli. We leverage a transformer architecture that can flexibly condition on a variable number of in-context image stimuli, learning an inductive bias over multiple subjects. During training, we explicitly optimize the model for in-context learning. By jointly conditioning on image features and voxel activations, our model learns to directly generate better performing voxelwise models of higher visual cortex. We demonstrate that BraInCoRL consistently outperforms existing voxelwise encoder designs in a low-data regime when evaluated on entirely novel images, while also exhibiting strong test-time scaling behavior. The model also generalizes to an entirely new visual fMRI dataset, which uses different subjects and fMRI data acquisition parameters. Further, BraInCoRL facilitates better interpretability of neural signals in higher visual cortex by attending to semantically relevant stimuli. Finally, we show that our framework enables interpretable mappings from natural language queries to voxel selectivity. Our code and model weights are publicly available at https://github.com/leomqyu/BraInCoRL.
@@ 1 Introduction
Human visual cortex transforms raw sensory input into behaviorally-relevant representations of the world. While early visual areas are characterized by retinotopic organization and selective tuning to simple features such as edges and orientation gradients [kuffler1953discharge, hubel1962receptive, atick1992does, verweij2003surround, grill2004human], higher-order visual areas demonstrate selectivity to more abstract semantics and categories. While this functional organization is largely consistent across individuals at a coarse scale, the spatial distribution and fine-grained semantic selectivity within visual cortex varies due to structural differences, developmental experience, and life-long learning [tarr2000ffa, gauthier2000expertise, willems2010cerebral, cai2013complementary, pinel2015genetic, saygin2016connectivity]. Such functional inter-subject differences pose a fundamental challenge in constructing generalizable models of higher visual cortex that can adapt to subject-specific neural organization without exhaustive data collection for every individual.
Recent advances in deep learning offer a promising avenue for addressing this challenge. Vision models pretrained on large-scale image datasets not only achieve strong object recognition performance, but also recapitulate hierarchical processing patterns observed in biological vision [yamins2014performance, wang2022incorporating, conwell2022can]. While these models may encapsulate some universal principles of visual processing [chen2024universal], they do not inherently account for individual differences in cortical organization. To close the gap between artificial and biological systems, researchers have developed image-computable fMRI encoders – models that predict brain activity from visual stimuli  [naselaris2011encoding]. These encoders typically regress image features onto voxelwise brain responses using subject-specific data, acting as computational probes of visual processing. Unfortunately, current approaches require many hours of costly fMRI scans per subject to fit these mappings – a prohibitive bottleneck for scalability to new populations, stimuli, and tasks, especially in clinical settings where collecting large amounts of data is difficult.
@@
Figure 1: BraInCoRL: Meta-Learning an In-Context Visual Cortex Encoder. (a) The voxelwise brain encoding problem. For each voxel, there is a response function that maps from visual stimuli to voxel activation. In practice, we can only observe the noisy measurements from fMRI. The goal is to infer an image-computable function for each voxel to predict its activation. (b) BraInCoRL treats each voxel as a meta-learning task, and samples (image, response) pairs from multiple subjects. During testing, the model is conditioned on a small number of novel images and measurements from a new subject and directly outputs the function parameters. (c) From left to right, the explained variance from the full model trained on 9,000 images from one subject, BraInCoRL with only 100100 in-context images from the new subject, and a baseline ridge regression also with 100100 images (for this baseline, voxelwise regularization is determined using 5-way cross-validation). Our method achieves much higher data efficiency than baseline. (d) Explained variance as a function of in‐context support set size. As the in-context support set size increases from 0 to 1,000, BraInCoRL steadily improves and approaches the fully trained reference model fit to converge on each subject’s full 9,000-image training set, demonstrating high prediction accuracy and data efficiency.
@@ 2 Related work
Computational modeling of neural data often involves two complementary approaches: encoding models that map from stimuli to neural activations, and decoding models that map from neural data to stimuli [naselaris2011encoding, kamitani2005decoding, norman2006beyond, han2019variational, seeliger2018generative, shen2019deep, ren2021reconstructing, dai2025mindalignerexplicitbrainfunctional, gifford2024opportunities]. The development of both approaches has been facilitated by advances in machine learning models. For encoding models, the dominant approach is to combine pretrained deep feature extractors with linear voxelwise weights [dumoulin2008population, gucclu2015deep, klindt2017neural, eickenberg2017seeing, wen2018neural, gaziv2022self]. More recent approaches have proposed to leverage transformers [adeli2023predicting, bao2025mindsimulator] to learn the relationship between brain regions of a single subject. Most similar to our framework is the pioneering work by Adeli et al. [adeli2023predicting] and Beliy & Wasserman et al. [beliy2024wisdom] which uses an auto-decoder based transformer network for multi-subject voxelwise encoding; However these approaches still require fine-tuning for novel subjects. More generally, encoders have been used to investigate the coding properties in higher-order visual areas [khosla2022high, khosla2022characterizing, efird2024s, yang2024brain, yang2024alignedcut, luo2024brain, Sarch2023.05.29.542635, lappe2024parallel]. Encoders have been further combined with image-generation models  [walker2019inception, bashivan2019neural, ponce2019evolving, ratan2021computational, gu2022neurogen, pierzchlewicz2023energy, luo2023brain, cerdas2024brainactiv] or language-generation models [luo2024brainscuba, matsuyama2025lavca] to explore semantic selectivity. Recent progress on large generative models has enabled stimulus reconstruction from fMRI, EEG, and MEG signals for images [takagi2022high, chen2023seeing, lu2023minddiffuser, ozcelik2023brain, doerig2022semantic, ferrante2023brain, liu2023brainclip, mai2023unibrain, scotti2024mindeye2, benchetrit2023brain, li2024visual, guo2024neuro3d3dvisualdecoding], videos [zhu2024multimodallatentvariablescrossindividual, schneider2023learnable, chen2023cinematic, gong2024neuroclips, yeung2024neural, liu2024eeg2video, fosco2024brain], and language/audio [pasley2012reconstructing, varoquaux2017assessing, bellier2023music, oota2023speech, jo2024eeg, willett2023high, metzger2023high].
@@
Human visual cortex exhibits a hierarchical organization from primary visual to higher-order visual areas. The higher visual cortex is characterized by a tiling of semantically specialization. Approaches using functional localizers have identified category selective regions in higher visual that are responsive to faces [sergent1992functional, allison1994human, mccarthy1997face, kanwisher1997fusiform], places [aguirre1996parahippocampus, epstein1998cortical], bodies [downing2001cortical], objects [grill2003neural, malach1995object], food [khosla2022, pennock2023color, Jain2023], and words [cohen2000visual, dehaene2001cerebral]. While the spatial location of these broad category-selective regions are generally consistent across subjects [op2008stable], significant inter-individual variability exists in their anatomical location, spatial extent, and response profiles [tarr2000ffa, gauthier2000expertise, willems2010cerebral, cai2013complementary, van2017development, golarai2015experience, liu2018successful, de2019factors, abbasi2020genetic]. Accurately characterizing visual processing in higher-order visual areas necessitates subject-specific encoding models that capture individual diversity [nieto2012subject].
@@
Our work builds upon meta-learning and in-context learning (ICL). Meta-learning trains models to ”learn to learn” from a distribution of tasks, enabling quick adaptation to new tasks with few examples, via methods like meta-optimization [finn2017model, nichol2018first, rajeswaran2019meta] or metric-based approaches [snell2017prototypical]. More recently, ICL has emerged as a powerful capability in Large Language Models [brown2020language, von2023transformers], where models adapt to new tasks at inference time solely based on examples provided in their prompt without any parameter updates [min2021metaicl, coda2023meta]. This has led to hypotheses that ICL is an emergent form of implicit meta-learning, where transformers effectively learn underlying learning algorithms during pre-training [garg2022can, dai2022can]. Our goal is to learn the structure of functions that map between visual stimuli and voxelwise brain response. Our framework combines meta-training (across voxels and subjects) and in-context learning (across stimuli) to enable training free adaptation to novel subjects.
@@ 3.2 Meta-Learning an In-Context Transformer
For a single voxel we define a support set of $p$ images and neural responses $\{(x_{1},\beta_{1}),(x_{2},\beta_{2}),...,(x_{p},\beta_{p})\}$, where $x_{i}\in\mathbb{R}^{m}$ is the image embedding vector extracted by a frozen image feature extractor $\phi(\cdot)$, i.e., $x_{i}=\phi(I_{i})$, and $\beta_{i}\in\mathbb{R}$ is the voxel’s response observed for image $I_{i}$. Each pair is concatenated to form context tokens $c_{i}=[x_{i};\beta_{i}]$, and the full context is defined as $\{c_{1},\dots,c_{p}\}$. Unlike traditional in-context inference in large language models, where there is a query concatenated to the end of the context, we avoid expensive per-voxel inference by directly generating the parameters for the voxelwise visual response function. During training, we optimize the BraInCoRL transformer $T$ with parameters $\theta$ such that it outputs voxel response function $f$ with parameters $\omega$:
@@ 3.3 Test-time Context Scaling
At test time, when we encounter a new subject, we assume we have access to a small set of novel images and the corresponding brain responses – we want to predict a voxelwise encoder. While our voxelwise parameterization successfully resolves the challenge of unknown test-time anatomy and geometry, the challenge of unknown test-time context size remains. Unlike transformers in language models, where the output is dependent on the order of the samples, we want our model to be explicitly invariant to input order. In order to facilitate variable length context, we utilize logit scaling [kexuefm-8823, chiang2022overcoming, bai2023qwen]. Assuming a query/key ($q,k$) with $d_{k}$ features and a length $l$ context:
$\alpha_{\text{orig}}=\frac{q\cdot k}{\sqrt{d_{k}}};\quad\alpha_{\text{scaled}}=\frac{\log{(l)}\cdot q\cdot k}{\sqrt{d_{k}}}$ (4)
We find this method effectively enables context scaling when trained with variable length context. A more detailed description on the test-time context scaling technique is provided in Appendix A.4. While the hypernetwork could, in principle, parameterize any class of neural encoders (e.g., MLPs, convolution, attention layers), prior studies utilizing brain data have largely used linear parameterizations that map from deep network features to voxel responses [wang2022incorporating, conwell2022can], and find that such a choice offers high performance and interpretability. Given features $x\in\mathbb{R}^{1\times q}$ from a pretrained neural network $x=\phi(\mathcal{I})$, we adopt the same setup and predict the final voxel response:
@@ 4 Experiments
Figure 3: Evaluation on NSD. (a) Prediction explained variance of BraInCoRL improves on novel subjects with larger in-context support set size, outperforming within-subject ridge regression and approaching the fully trained reference model fit on each subject’s full 9,000-image training set, using far less data. (b) Ablation (100 support images) comparing BraInCoRL variants: the original model trained while holding out the novel subject’s 9,000 test-time support images (“HO”), a BraInCoRL model trained without this holdout (“no HO”), and a pretraining-only BraInCoRL model, alongside the within-subject ridge baseline. Results show that finetuning with real fMRI data improves performance, and holding out the test subject’s image data does not hinder generalization. (c) Voxelwise explained variance from BraInCoRL (100 images) is strongly correlated with fully trained reference models across different visual encoder backbones. Note that the y-axis represents explained variance of the fully trained model (9,000 images), while x-axis represents explained variance of BraInCoRL.
Table 1: Voxelwise performance across five category‐selective regions. Explained variance is shown for our in‐context model (“BraInCoRL”) that uses just 100 in-context images, the fully trained reference model fit to converge on each subject’s full 9,000-image training set (“Fully Trained”), and within-subject ridge regression baselines (100, 300 within-subject test images), plus the FsAverage map averages over other subjects. Our model outperforms both subject-wise and anatomical baselines, and demonstrates strong data-efficiency.
@@ 4.1 Setup
We primarily perform experiments with the Natural Scenes Dataset (NSD) [allen2022massive], but then validate with the BOLD5000 dataset [chang2019bold5000]. Both are large-scale neural datasets: NSD is the largest 7T fMRI image viewing dataset available, where eight subjects each viewed $\sim 10,000$ images; BOLD5000 is a 3T dataset, where four subjects each viewed $\sim 5000$ images. In NSD each image was viewed up to three times, while in BOLD5000 only a subset of images were viewed up to four times. For NSD, of the eight subjects, four subjects (S1, S2, S5, S7) completed scanning and are the focus of our analysis in the main paper. The results of other subjects are presented in the supplemental materials. For each subject, $\sim 9,000$ images are unique to each other, while $\sim 1,000$ are viewed by all eight subjects. The NSD stimuli were sourced from the MS COCO dataset (as were a subset of the BOLD5000 stimuli). Unless otherwise noted, we perform our analysis in subject native volume space (func1pt8mm) for NSD, where the voxelwise betas are z-scored within each session then averaged across repeats of the same stimulus. In order to rigorously evaluate BraInCoRL for a given subject, we use the 3 × 9,000 unique images viewed by the other three subjects as the meta-training data. During the ROI-wise evaluation for NSD, we follow prior work [luo2023brain] and apply a $t$-statistic cutoff of $t>2$ by leveraging independent functional localizer data provided with the dataset to threshold the originally broad definitions. During quantitative evaluation, we follow prior work [conwell2024large] and apply a voxel quality cutoff of $\texttt{ncsnr}>0.2$. For BOLD5000, we use a model trained on four NSD subjects (S1, S2, S5, S7). Following the suggestion of BOLD5000 authors, we only model stimuli with 4 repeats and apply a cutoff of $\texttt{ncsnr}>0.3$. Voxel responses are averaged across repeats.